faithkristine:

the airplane door is 19 feet off the ground and the ramp has a 31 degree angle of evaluation. what is the length y of the ramp?

Eiwoh2:

@AZ

AZ:

Let's draw the image, okay? 

faithkristine:

ok

AZ:

Does that make sense? The ramp is on the ground and at an angle of 31 degrees to reach the door which is 19 feet high and we want to calculate the length of the ramp

faithkristine:

yeah but how

AZ:

Have you learned about sin cos or tan?

faithkristine:

yes but i dont get it

AZ:

Sure, we can cover it again

AZ:

Have you ever heard of SOH CAH TOA Essentially, it's an acronym to help you remember the trigonometric functions sin of an angle equals the opposite side over the hypotenuse We write that as sin \(\theta\) = opposite/hypotenuse Do you see the initials? Sin = Opposite/ Hypotenuse SOH Similarly, for cosine and tangent cos \(\theta\) = adjacent/hypotenuse CAH tan \(\theta\) = opposite/adjacent TAH

faithkristine:

are you a teacher

AZ:

Here's an image to help you see it together

1 attachment
faithkristine:

ok

AZ:

No haha, I'm just a student like you :)

faithkristine:

cool:)

AZ:

So in the image I drew, do you see that \(\theta\) is going to be the angle which is 31 We're looking for the length of the ramp which is the hypotenuse And we know the side opposite of the angle So would we use sin, cos, or tan?

faithkristine:

ok

faithkristine:

im sorry but its late can you just tell me the answer and we can go over it tomorrow

AZ:

I would love to but the purpose of this site is to help teach and learn. If I gave you the answer, I'd be breaking the rules and get in trouble It'll only take like 3 minutes if you work with me to get to the correct answer!

faithkristine:

ok

AZ:

Remember sin = opposite / hypotenuse cos = adjacent / hypotenuse tan = opposite / adjacent we know the side OPPOSITE to the angle and we want to know the hypotenuse which one do we use? which one has both the opposite and the hypotenuse

faithkristine:

sin?

AZ:

Exactly! So remember how we set it up? \(\sin \theta = \dfrac{\text{opposite}}{\text{hypotenuse}}\) so our angle or \(\theta\) is 31 and the side opposite is 19 do you know how to solve for the hypotenuse? (You have to use a calculator to find sin 19 by the way)

faithkristine:

no how do you put into calculator

AZ:

You can use Google but you have to be careful that you're doing sin 31 degrees otherwise it'll do 31 radians which is not what we want so https://www.google.com/search?q=sin+31+degrees

AZ:

So what is sin 31 degrees = ?

faithkristine:

0.5150

faithkristine:

??

AZ:

Good! So now we have \(\sin \theta = \dfrac{\text{opposite}}{\text{hypotenuse}}\) \(\sin 31^o = \dfrac{19}{\text{hypotenuse}}\) \( 0.5150 = \dfrac{19}{\text{hypotenuse}}\) and think of the hypotenuse like it's an 'x' \( 0.5150 = \dfrac{19}{x}\) How would you solve for x?

faithkristine:

why do i need t solve for x its just y

faithkristine:

oh wait nvm haha

faithkristine:

im so dumb

faithkristine:

i dont know how to solve for x

AZ:

haha yeah, we can use any letter but 'x' is the one that's most commonly used in math

faithkristine:

so 0.52 is the answer cause you round it to the tenth place

AZ:

That's not a problem \( 0.5150 = \dfrac{19}{x}\) So we have 'x' in the denominator. What happens if we multiply 'x' on both sides So essentially, how would you simplify \( 0.5150 \times x= \dfrac{19}{x} \times x\)

faithkristine:

ok so how would you simplify it

AZ:

@faithkristine wrote:
so 0.52 is the answer cause you round it to the tenth place
No no, that's not the final answer. That was just the value of sin 31 degrees Also, think about it. Does it make sense for the ramp to only be 0.52 feet long when it's 19 feet tall?

faithkristine:

true

AZ:

@faithkristine wrote:
ok so how would you simplify it
Do you see how the 2 x on the left side cancel out since one is in the numerator and the other is in the denominator?

faithkristine:

yeah

AZ:

So we're left with 0.5150 * x = 19 can you solve for x now?

faithkristine:

no how do you put that in a calculator

AZ:

uhh you wouldn't be able to put it into a normal calculator but this is something you should be able to do on your own now

faithkristine:

im dumb trust me i cant

AZ:

you want to get 'x' all by itself right now, it's being multiplied with 0.5150 so to get x all by itself, and essentially move the 0.5150 to the other side, you have to do the opposite of multiplication what is the opposite of multiplication?

faithkristine:

division

AZ:

Exactly!! So we divide 0.5150 on both sides So all you need to do is divide 19 by 0.5150 x = 19 / 0.5150 x = ??

faithkristine:

36.9

AZ:

There you go!

faithkristine:

thank you so much ill keep in touch

AZ:

Of course! You're welcome!